OpenStudy (anonymous):

What is the reciprocal of 8? -8 -1/8 8/1 1/8

OpenStudy (lgbasallote):

hint reciprocal of 2 is \(\frac 12\) reciprocal of 3 is \(\frac 13\) reciprocal of x is \(\frac 1x\) does that give you any idea?

OpenStudy (anonymous):

oh so is it 1/8

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

thanks :)

OpenStudy (lgbasallote):

welcome